OpenStudy (anonymous):

Integrate: (2xˆ2+3)/xˆ2

sam (.sam.):

Same that I've thought you, \[\int\limits \frac{2 x^2+3}{x^2} \, dx\] Expand out fraction \[\int\limits \frac{3}{x^2}+2 \, dx\] \[\int\limits\limits 3x^{-2}+2 \, dx\] Integrate using power rule